THE SUBSTATION BUSBAR IS MADE UP OF 2 INCHES ROUND COPPER BARS 240in LONG.




WHAT IS THE RESISTANCE OF EACH BAR IF RESISTIVITY IS 1.724 x 10-6 OHM-CM?

1
Expert's answer
2021-11-16T13:49:25-0500

R=2in="0.0254\times2=0.0508m"

L=240 in "=240\times0.0254=6.096m"

"\rho=1.724\times10^{-8}\Omega m"

We know that

"R=\frac{\rho L}{A}"

Put value


"R=\frac{1.724\times10^{-8}\times6.096}{3.14\times0.0508^2}=1.296\times10^{-05}\Omega"